Circular No. 8177

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

  URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF (22) KALLIOPE AND (45) EUGENIA As noted on MPC 49221, the responsibility in the IAU for naming

satellites of minor planets was recently transferred from the Working Group on Planetary System Nomenclature (WGPSN) to the Committee on Small Bodies Nomenclature (CSBN), which at its meetings in Sydney in July adopted the name “Linus” for the satellite (22) Kalliope I = S/2001 (22) 1 (cf. IAUC 7703). In the same vein, three years ago, the WGPSN adopted the name “Petit Prince” for the satellite (45) Eugenia I = S/1998 (45) 1 (cf. IAUC 7129).

SATELLITES OF JUPITER, SATURN, URANUS The IAU WGPSN has approved the following new designations and

names of satellites of major planets (cf. IAUC 5347, 7479, 7998, 8023):

Jupiter XXVIII Autonoe = S/2001 J 1
Jupiter XXIX Thyone = S/2001 J 2
Jupiter XXX  Hermippe = S/2001 J 3
Jupiter XXXI Aitne = S/2001 J 11
Jupiter XXXII  Eurydome = S/2001 J 4
Jupiter XXXIII Euanthe  = S/2001 J 7
Jupiter XXXIV Euporie  = S/2001 J 10
Jupiter XXXV  Orthosie = S/2001 J 9
Jupiter XXXVI Sponde  = S/2001 J 5
Jupiter XXXVII Kale  = S/2001 J 8
Jupiter XXXVIII Pasithee  = S/2001 J 6
Saturn XIX  Ymir  = S/2000 S 1
Saturn XX  Paaliaq  = S/2000 S 2
Saturn XXI  Tarvos  = S/2000 S 4
Saturn XXII  Ijiraq  = S/2000 S 6
Saturn XXIII  Suttung  = S/2000 S 12
Saturn XXIV  Kiviuq  = S/2000 S 5
Saturn XXV  Mundilfari = S/2000 S 9
Saturn XXVI  Albiorix  = S/2000 S 11
Saturn XXVII  Skadi  = S/2000 S 8
Saturn XXVIII Erriapo  = S/2000 S 10
Saturn XXIX  Siarnaq  = S/2000 S 3
Saturn XXX  Thrym  = S/2000 S 7
Uranus XXI  Trinculo  = S/2001 U 1
2003 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green